If a convex lens is cut along diameter into two equal parts,what will be the focal length of each part..
1
Expert's answer
2021-08-19T14:24:55-0400

The focal length for each turns out to be double the original focal length. You can solve this from combination lens equation:

"\large\frac{1}{f} = \frac{1}{f_1} + \frac{1}{f_2}"

Where the original focal length is f, and the focal length of each half is "f_1" and "f_2"

Since the two halves are identical, then "f_1 = f_2" . So the equation becomes:

"\large\frac{1}{f} = \frac{1}{f_1} + \frac{1}{f_2} = \frac{2}{f_1} \space" "or \space \space f = \large\frac{f_1}{2}"

Solving for one of the lens half:  "f_1 = 2f."